A beam of length 10m is supported at its ends and weighs 100kg/m. The ODE governing Cy" = 100{(x^2)/2 - 5x) where C is a constant. Solve this equation to determine the deflection at x=6m , given that y=0 at x=0 and x=10.
1
Expert's answer
2020-06-11T20:41:30-0400

"Cy" = 100(\frac {x^2}{2} - 5x)"


integrating,


"Cy'= 100(\frac {x^3}{6} - 5\frac {x^2}{2})+c_{1}"


again Integrating,


"Cy= 100(\frac {x^4}{24} - 5\frac {x^3}{6})+c_{1} x+c_{2}"


at y=0 at x=0 "c_{2}=0"


and y =0 and x = 10 "c_1=4.17 \times 10^3"


deflection at x=6 m


"y= \frac {12420} {C}"